# algorithmic

### v4

@@ Abstract
Recently, the equivariance of models with respect to a group action has become an important topic of research in machine learning. Analysis of the built-in equivariance of existing neural network architectures, as well as the study of methods for building model architectures that explicitly “bake in” equivariance, have become significant research areas in their own right. However, imbuing an architecture with a specific group equivariance imposes a strong prior on the types of data transformations that the model expects to see. While strictly-equivariant models enforce symmetries, such as those due to rotations or translations, real-world data does not always conform to such strict equivariances, be it due to noise in the data or underlying physical laws that encode only approximate or partial symmetries. In such cases, the prior of strict equivariance can actually prove too strong and cause models to underperform on real-world data. Therefore, in this work we study a closely related topic, that of almost equivariance. We provide a definition of almost equivariance that differs from those extant in the current literature and give a practical method for encoding almost equivariance in models by appealing to the Lie algebra of a Lie group. Specifically, we define Lie algebra convolutions and demonstrate that they offer several benefits over Lie group convolutions, including being well-defined for non-compact Lie groups having non-surjective exponential map. From there, we pivot to the realm of theory and demonstrate connections between the notions of equivariance and isometry and those of almost equivariance and almost isometry. We prove two existence theorems, one showing the existence of almost isometries within bounded distance of isometries of a general manifold, and another showing the converse for Hilbert spaces. We then extend these theorems to prove the existence of almost equivariant manifold embeddings within bounded distance of fully equivariant embedding functions, subject to certain constraints on the group action and the function class. Finally, we demonstrate the validity of our approach by benchmarking against datasets in fully equivariant and almost equivariant settings.
@@ 1 Introduction
The past few years have shown a surge in interest in equivariant model architectures, those that explicitly impose symmetry with respect to a particular group acting on the model inputs. While data augmentation strategies have been proposed to make generic models exhibit greater symmetry without the need for equivariant model architectures, much work has demonstrated that this is an inefficient approach at best (Gerken et al., 2022b; Lafarge et al., 2020; Wang et al., 2022b). As such, developing methods for building neural network layers that are equivariant to general group actions is of great importance.
@@ 2.1 Strict Equivariance
Much of the work in developing strictly-equivariant model architectures began with the seminal paper of Cohen & Welling (2016), which introduced the group equivariant convolutional neural network layer. Kondor & Trivedi (2018) generalized this notion of equivariance and convolution to the action of an arbitrary compact group. Further generalizations followed, with the creation of convolutions (Finzi et al., 2020) and efficient MLP layers (Finzi et al., 2021a) equivariant to arbitrary Lie groups. Other neural network types have also been studied through the lens of equivariance, for example, graph neural networks (Satorras et al., 2021), (Batzner et al., 2022), transformers (Hutchinson et al., 2021a), and graph transformers (Liao & Smidt, 2023). Cohen et al. (2019) consolidated much of this work into a general framework via which equivariant layers can be understood as maps between spaces of sections of vector bundles. Similar to our work, Dehmamy et al. (2021) devised a convolutional layer on the Lie algebra designed to approximate group convolutional layers. However, their objective was to make the layer as close to equivariant as possible whereas our layer is designed to be flexible so as to be capable of modelling almost equivariances. Finally, rather than devising a new equivariant layer type, Gruver et al. (2023) developed a method based on the Lie derivative which can be used to detect the degree of equivariance learned by an arbitrary model architecture.
@@ 2.2 Almost Equivariance
One of the first works on almost equivariance was Finzi et al. (2021b), which introduced the Residual Pathway Prior. Their idea is to construct a neural network layer, $f$, that is the sum of two components, $A$ and $B$, where $A$ is a strictly equivariant layer and $B$ is a more flexible, non-equivariant layer
@@
The approach taken in Wang et al. (2022a) is somewhat different. They give an explicit definition of approximate equivariance, then model it via a relaxed group convolutional layer wherein the single kernel, $\Psi$, of a strictly equivariant group convolutional layer is replaced with a set of kernels, ${\{\Psi_{l}\}}_{l=1}^{L}$. This introduces a specific, symmetry-breaking dependence on a pair of group elements, $(g,h)$, i.e.
@@
Romero & Lohit (2022) take an altogether different approach. They introduce a model, which they call the Partial G-CNN, and show how to train it to learn layer-wise levels of equivariance from data. A key differentiator in their method is the learning of a probability distribution over group elements at each group convolutional layer, allowing them to sample group elements during group convolutions.
@@
where $p(u)$ is a probability distribution on $G$ and $\mu_{G}$ is the Haar measure.
@@
van der Ouderaa et al. (2022) relax equivariance constraints by defining a non-stationary group convolution
@@
Finally, Petrache & Trivedi (2023) provide a take on approximate equivariance rooted in statistical learning theory and provide generalization and error bounds on approximately equivariant architectures.
@@ 2.3 Notions of Approximate Equivariance
The first type occurs when we only have partially observed data, for example, a single pose of a 3D object captured in a 2D image or an object occlusion in computer vision. Wang et al. (2023) refer to this as extrinsic equivariance in that applying a group transformation to an in-distribution data point transforms it to an out-of-distribution data point. This type of partial equivariance is often addressed via data augmentation. We do not explicitly test our approach in this setting.
@@
The third type occurs when we have data that naturally exhibits almost equivariance. For example, data sampled from vector fields and PDEs governing natural physical processes often exhibit this quality. This is another setting we exxplicitly address.
Finally, there is what Wang et al. (2023) call incorrect equivariance. This occurs when applying a group transformation to a data point qualitatively and quantitatively changes its label. For example, rotating the digit 6 by 180 degrees turns it into the digit 9 and vice versa. We do not explicitly address this in our method, but our model performs competitively on the Rot-MNIST classification task, indicating that it has the capability of accounting for incorrect equivariances in its modeling.
@@ 3.1 Equivariance & Almost Equivariance
Definition 3.2 ($\varepsilon$-almost equivariant function).
@@
This definition is reminiscent of one given in a question posed by Stanislaw Ulam (Ulam, 1960) concerning the stability of certain “quasi” group homomorphisms. In particular, given a group $\Gamma$, a group $G$ equipped with a distance $d$, and a $\delta$-homomorphism, $\mu:\Gamma\to G$, satisfying
@@ 3.2 Isometries, Isometry Groups, and Almost Isometries
We begin our discussion of the theory underlying almost equivariance by studying the notions of isometry and almost isometry. Because we often seek to impose in our models equivariance with respect to the action of the isometry group of a manifold from which data is sampled, we find it worthwhile to study isometries as a precursor to studying equivariance. An isometry is a mapping of metric spaces that preserves the distance metric. Some common types of metric spaces for which there exists a natural notion of isometry are normed spaces, such as Banach spaces, and Riemannian manifolds. In this work, we focus most of our analysis on Riemannian manifolds, as they are among the most general spaces upon which equivariant models operate and underlie all of geometric deep learning (Bronstein et al., 2021).
Definition 3.3 (isometry of a Riemannian manifold).
Let $(M,g)$ and $(\tilde{M},\tilde{g})$ be Riemannian manifolds, $\varepsilon>0$, and $\varphi:M\to\tilde{M}$ a diffeomorphism. Then we say $\varphi$ is an isometry of $M$ if $g=\varphi^{*}\tilde{g}$. In other words, the metric $\tilde{g}$ can be pulled back by $\varphi$ to get the metric $g$.
@@
Definition 3.4 ($\varepsilon$-almost isometry of a Riemannian manifold – local version).
@@
Definition 3.5 ($E$-almost isometry of a Riemannian manifold – global version).
Given Riemannian manifolds $(M,g)$ and $(M,\tilde{g})$ and a local $\varepsilon$-almost isometry, $\varphi:M\to M$, we say that $\varphi$ is a global $E$-almost isometry if there exists a scalar field, $E:M\to\mathbb{R}$, such that for any vector fields $X,Y\in\Gamma(TM)$, we have
@@
In particular, $\sum_{p\in M}E_{p}\leq\varepsilon\cdot\text{Vol}_{g}(M)$.
@@
Definition 3.6 (Isometry group of a Riemannian manifold).
@@
Theorem 3.7 (Myers & Steenrod (1939)).
@@
Unfortunately, it is relatively rare for an isometry group of a manifold to be abelian. Furthermore, we cannot, without some work, consider this theorem in the context of $\text{Iso}_{\varepsilon}(M)$ because the set of $\varepsilon$-almost isometries of a manifold does not form a group. To see why, note that composing two $\varepsilon$-almost isometries produces, in general, a $2\varepsilon$-almost isometry, thus the set of $\varepsilon$-almost isometries of a manifold is not closed under composition. Still, we can impose the abelian condition on group actions as a stepping stone towards studying more general group actions, almost isometries, and equivariant functions. Under the assumption of an abelian Lie group acting on a Riemannian manifold, we prove the following theorem.
@@
[…]
@@
$d(g\cdot f(x),f_{\varepsilon}(g\cdot x))$ $<d(g\cdot f(x),f(g\cdot x))+d(f(g\cdot x),f_{\varepsilon}(g\cdot x))$ (1)
@@
$d(g\cdot f_{\varepsilon}(x),f_{\varepsilon}(g\cdot x))$ $<d(g\cdot f_{\varepsilon}(x),g\cdot f(x))+d(g\cdot f(x),f_{\varepsilon}(g\cdot x))$ (3)
@@
Under the stated assumptions of $M$ a compact Riemannian manifold and $G$ a compact Lie group acting on $M$ by isometries, we can get the existence of $\rho$ and $\varphi$ by invoking the main theorem of Moore & Schlafly (1980). From there, note that setting $f_{\varepsilon}=\varphi(f)$ trivially satisfies (1) for any $\varepsilon$, although we seek a non-trivial solution. We can choose an arbitrary $x_{0}\in M$, and define $f_{\varepsilon}(x)=\varphi(f(x))$ for all $x\neq x_{0}\in M$. Next, since $M$ is compact, $\varphi(f)$ is bounded on $M$. We can then take a neighborhood $U$ of $x_{0}$ such that $B_{\varepsilon}(\varphi(f(x_{0})))\subseteq\varphi(U)$. We can then choose an arbitrary $y\in B_{\varepsilon}(\varphi(f(x_{0})))$, while requiring $y\neq\varphi(f(x_{0}))$, and set $f_{\varepsilon}(x_{0})=y$. Then $f_{\varepsilon}$ is an $\varepsilon$-almost isometric embedding of $M$, but $f_{\varepsilon}\neq f$, as desired. Furthermore, given a suitable topology on $C^{\infty}(M)$ (such as the compact-open topology), $B_{\varepsilon}(\varphi(f(x_{0})))$ is open so that there exist infinitely many such $f_{\varepsilon}\neq f$, and they can be taken to be continuous. ∎
@@
3.3 Ulam Stability Theory & Fickett’s Theorem
@@
(Hyers & Ulam (1945)) Let $E$ be a complete real Hilbert space. Let $\varepsilon>0$ and $T$ be a surjection of E into itself that is an $\varepsilon$-isometry, that is, $|\rho(T(x),T(y))-\rho(x,y)|<\varepsilon$, for all $x,y\in E$, where $\rho$ denotes the inner product in $E$. Assume that $T(0)=0$. Then the limit
@@
Taken together, these two results show that no matter what $\varepsilon$-almost isometry we define, it is never “far off” from a full isometry, with the distance between the two bounded above by $27\varepsilon^{1/2^{n}}$. Most recently, Väisälä (2002) proved an even tighter bound, but its discussion is beyond the scope of this paper.
To apply Theorem 3.11 and Theorem 3.12 in the context of machine learning, note that by the Nash Embedding Theorem (Nash, 1954), we can smoothly and isometrically embed any Riemannian manifold $(M,g)$ into $\mathbb{R}^{n}$ for some $n$. If $M$ is compact, then the embedding of $M$ in $\mathbb{R}^{n}$ will be a compact, and therefore bounded, subset of $\mathbb{R}^{n}$. We can then apply Theorem 3.12 or Theorem 3.11 to any $\varepsilon$-isometry of $M$ to get a nearby isometry of $M$ as a subset of $\mathbb{R}^{n}$.
If $M$ is not compact, let $S\subseteq\mathbb{R}^{n}$ be its smooth isometric embedding. We can then apply Theorem 11.4 of Wells & Williams (1975), which states that for a finite-dimensional Hilbert space, $H$, we can extend any isometry of $S$ to an isometry on the linear span of $S$. Assuming the completion, $\bar{S}$, of $S$ is contained in the linear span of $S$, we can then, for any surjective $\varepsilon$-isometry of $\bar{S}$ into itself, apply Theorem 3.11 to recover an isometry of $\bar{S}$.

### v6

# Almost Equivariance via Lie Algebra Convolutions

@@ Abstract
Recently, equivariant neural networks have become an important topic of research in machine learning. However, imbuing an architecture with a specific group equivariance imposes a strong prior on the types of data transformations that the model expects to see. While strictly-equivariant models enforce symmetries, real-world data does not always conform to such strict equivariances. In such cases, the prior of strict equivariance can actually prove too strong and cause models to underfit on real-world data. Therefore, in this work we study a closely related topic, that of almost equivariance. We provide a definition of almost equivariance that differs from those extant in the current literature and give a practical method for encoding almost equivariance in models by appealing to the Lie algebra of a Lie group. Specifically, we define Lie algebra convolutions and demonstrate that they offer several benefits over Lie group convolutions, including being well-defined for non-compact Lie groups having non-surjective exponential map. From there, we pivot to the realm of theory and demonstrate parallel connections between the notions of equivariance and isometry and those of almost equivariance and almost isometry. Finally, we demonstrate the validity of our approach by benchmarking against datasets in fully equivariant and almost equivariant settings.
@@ 1 Introduction
The past few years have shown a surge in interest in equivariant model architectures, those that explicitly impose symmetry with respect to a particular group acting on the model inputs. While data augmentation strategies have been proposed to make generic models exhibit greater symmetry without the need for equivariant model architectures, much work has demonstrated that this is an inefficient approach at best [15, 26, 45]. As such, developing methods for building neural network layers that are equivariant to general group actions is of great importance.
@@ 2.1 Strict Equivariance
Much of the work in developing strictly-equivariant model architectures began with the seminal paper of Cohen and Welling [4], which introduced the group equivariant convolutional neural network layer. Kondor and Trivedi [25] generalized this notion of equivariance and convolution to the action of an arbitrary compact group. Further generalizations followed, with the creation of convolutions [10] and efficient MLP layers [11] equivariant to arbitrary Lie groups. Other neural network types have also been studied through the lens of equivariance, for example, graph neural networks [39], [1], transformers [20], and graph transformers [30]. Cohen et al. [5] consolidated much of this work into a general framework via which equivariant layers can be understood as maps between spaces of sections of vector bundles. Similar to our work, Dehmamy et al. [6] devised a convolutional layer on the Lie algebra designed to approximate group convolutional layers. However, their objective was to make the layer as close to equivariant as possible whereas our layer is designed to be flexible so as to be capable of modelling almost equivariances. Finally, rather than devising a new equivariant layer type, Gruver et al. [16] developed a method based on the Lie derivative which can be used to detect the degree of equivariance learned by an arbitrary model architecture.
@@ 2.2 Almost Equivariance
One of the first works on almost equivariance was Finzi et al. [12], which introduced the Residual Pathway Prior. Their idea is to construct a neural network layer, $f$, that is the sum of two components, $A$ and $B$, where $A$ is a strictly equivariant layer and $B$ is a more flexible, non-equivariant layer
@@
The approach taken in Wang et al. [44] is somewhat different. They give an explicit definition of approximate equivariance, then model it via a relaxed group convolutional layer wherein the single kernel, $\Psi$, of a strictly equivariant group convolutional layer is replaced with a set of kernels, ${\{\Psi_{l}\}}_{l=1}^{L}$. This introduces a specific, symmetry-breaking dependence on a pair of group elements, $(g,h)$, i.e.
@@
Romero and Lohit [38] take an altogether different approach. They introduce a model, which they call the Partial G-CNN, and show how to train it to learn layer-wise levels of equivariance from data. A key differentiator in their method is the learning of a probability distribution over group elements at each group convolutional layer, allowing them to sample group elements during group convolutions.
@@
for $u\in S^{2},v\in S^{1}$, where $p(u)$ is a probability distribution on $G$ and $\mu_{G}$ is the Haar measure.
@@
van der Ouderaa et al. [42] relax equivariance constraints by defining a non-stationary group convolution
@@
Finally, Petrache and Trivedi [36] provide a take on approximate equivariance rooted in statistical learning theory and provide generalization and error bounds on approximately equivariant architectures.
@@ 2.3 Notions of Approximate Equivariance
The first type occurs when we only have partially observed data, for example, a single pose of a 3D object captured in a 2D image or an object occlusion in computer vision. Wang et al. [43] refer to this as extrinsic equivariance in that applying a group transformation to an in-distribution data point transforms it to an out-of-distribution data point. This type of partial equivariance is often addressed via data augmentation. We do not explicitly test our approach in this setting.
@@
The third type occurs when we have data that naturally exhibits almost equivariance. For example, data sampled from vector fields and PDEs governing natural physical processes often exhibit this quality. This is another setting we explicitly address.
Finally, there is what Wang et al. [43] call incorrect equivariance. This occurs when applying a group transformation to a data point qualitatively and quantitatively changes its label. For example, rotating the digit 6 by 180 degrees turns it into the digit 9 and vice versa. We do not explicitly address this in our method, but our model performs competitively on the Rot-MNIST classification task, indicating that it has the capability of accounting for incorrect equivariances in its modeling.
@@ 3.1 Equivariance & Almost Equivariance
Definition 3.2.
@@
This definition is reminiscent of one given in a question posed by Stanislaw Ulam [40] concerning the stability of certain “quasi” group homomorphisms. In particular, given a group $\Gamma$, a group $G$ equipped with a distance $d$, and a $\delta$-homomorphism, $\mu:\Gamma\to G$, satisfying
@@ 3.2 Isometries, Isometry Groups, and Almost Isometries
We begin our discussion of the theory underlying almost equivariance by studying the notions of isometry and almost isometry. Because we often seek to impose in our models equivariance with respect to the action of the isometry group of a manifold from which data is sampled, we find it worthwhile to study isometries as a precursor to studying equivariance. An isometry is a mapping of metric spaces that preserves the distance metric. Some common types of metric spaces for which there exists a natural notion of isometry are normed spaces, such as Banach spaces, and Riemannian manifolds. In this work, we focus most of our analysis on Riemannian manifolds, as they are among the most general spaces upon which equivariant models operate [3].
Definition 3.3.
Let $(M,g)$ and $(\tilde{M},\tilde{g})$ be Riemannian manifolds adding $\varphi:M\to\tilde{M}$ a diffeomorphism. Then we say $\varphi$ is an isometry of $M$ if $g=\varphi^{*}\tilde{g}$. In other words, the metric $\tilde{g}$ can be pulled back by $\varphi$ to get the metric $g$.
@@
Definition 3.4.
@@
Definition 3.5.
Given oriented and compact Riemannian manifolds $(M,g)$ and $(M,\tilde{g})$ and a local $\varepsilon$-almost isometry, $\varphi:M\to M$, we say that $\varphi$ is a global $E$-almost isometry if there exists a continuous, compactly-supported scalar field, $E:M\to\mathbb{R}$, such that for any normalized vector fields $X,Y\in\Gamma(TM)$, we have
@@
In particular, $\int_{M}E\ \omega_{g}\leq\varepsilon\cdot\text{Vol}_{g}(M)$, where $\omega_{g}$ is the canonical Riemannian volume form on $(M,g)$.
@@
Definition 3.6.
@@
Theorem 3.7 (Myers and Steenrod [34]).
@@
However, we cannot, without some work, consider this theorem in the context of $\text{Iso}_{\varepsilon}(M)$ because the set of $\varepsilon$-almost isometries of a manifold does not form a group. To see why, note that composing two $\varepsilon$-almost isometries produces, in general, a $2\varepsilon$-almost isometry, thus the set of $\varepsilon$-almost isometries of a manifold is not closed under composition. Still, we can impose the abelian condition on group actions as a stepping stone towards studying more general group actions, almost isometries, and equivariant functions. Under the assumption of an abelian Lie group acting on a Riemannian manifold, we prove the following theorem.
@@
Let $(M,g)$ be a Riemannian manifold and suppose its group of isometries, $G=\text{Iso}(M)$, is an abelian Lie group. Let $f\in\text{Iso}(M)$, and suppose there exists a continuous $\varepsilon$-almost isometry, $f_{\varepsilon}\in\text{Iso}_{\varepsilon}(M)$, with $f\neq f_{\varepsilon}$, such that
@@
$d(g\cdot f(x),f_{\varepsilon}(g\cdot x))$ $<d(g\cdot f(x),f(g\cdot x))+$ (1)
$d(f(g\cdot x),f_{\varepsilon}(g\cdot x))$ (2)
@@
$d(g\cdot f_{\varepsilon}(x),f_{\varepsilon}(g\cdot x))$ $<d(g\cdot f_{\varepsilon}(x),g\cdot f(x))+$ (4)
$d(g\cdot f(x),f_{\varepsilon}(g\cdot x))$ (5)
@@
Under the stated assumptions of $M$ a compact Riemannian manifold and $G$ a compact Lie group acting on $M$ by isometries, we can get the existence of $\rho$ and $\varphi$ by invoking the main theorem of Moore and Schlafly [33]. From there, note that setting $f_{\varepsilon}=\varphi(f)$ trivially satisfies (1) for any $\varepsilon$, although we seek a non-trivial solution. We can choose an arbitrary $x_{0}\in M$, and define $f_{\varepsilon}(x)=\varphi(f(x))$ for all $x\neq x_{0}\in M$. Next, since $M$ is compact, $\varphi(f)$ is bounded on $M$. We can then take a neighborhood $U$ of $x_{0}$ such that $B_{\varepsilon}(\varphi(f(x_{0})))\subseteq\varphi(U)$. We can then choose an arbitrary $y\in B_{\varepsilon}(\varphi(f(x_{0})))$, while requiring $y\neq\varphi(f(x_{0}))$, and set $f_{\varepsilon}(x_{0})=y$. Then $f_{\varepsilon}$ is an $\varepsilon$-almost isometric embedding of $M$, but $f_{\varepsilon}\neq f$, as desired. Furthermore, given a suitable topology on $C^{\infty}(M)$ (such as the compact-open topology), $B_{\varepsilon}(\varphi(f(x_{0})))$ is open so that there exist infinitely many such $f_{\varepsilon}\neq f$, and they can be taken to be continuous. ∎
@@
3.3 Ulam Stability Theory
@@
(Hyers and Ulam [22]) Let $E$ be a complete real Hilbert space. Let $\varepsilon>0$ and $T$ be a surjection of E into itself that is an $\varepsilon$-isometry, that is, $|\rho(T(x),T(y))-\rho(x,y)|<\varepsilon$, for all $x,y\in E$, where $\rho$ denotes the inner product in $E$. Assume that $T(0)=0$. Then the limit
@@
Taken together, these two results show that no matter what $\varepsilon$-almost isometry we define, it is never “far off” from a full isometry, with the distance between the two bounded above by $27\varepsilon^{1/2^{n}}$. Most recently, Väisälä [41] proved an even tighter bound, but its discussion is beyond the scope of this paper.
To apply Theorem 3.11 in the context of machine learning, note that by the Nash Embedding Theorem [35], we can smoothly and isometrically embed any Riemannian manifold $(M,g)$ into $\mathbb{R}^{n}$ for some $n$. If $M$ is compact, then the embedding of $M$ in $\mathbb{R}^{n}$ will be a compact, and therefore bounded, subset of $\mathbb{R}^{n}$. We can then apply Theorem 3.11 to any $\varepsilon$-isometry of $M$ to get a nearby isometry of $M$ as a subset of $\mathbb{R}^{n}$.
If $M$ is not compact, let $S\subseteq\mathbb{R}^{n}$ be its smooth isometric embedding. We can then apply Theorem 11.4 of Wells and Williams [48], which states that for a finite-dimensional Hilbert space, $H$, we can extend any isometry of $S$ to an isometry on the linear span of $S$. Assuming the completion, $\bar{S}$, of $S$ is contained in the linear span of $S$, we can then, for any surjective $\varepsilon$-isometry of $\bar{S}$ into itself, apply Theorem 3.11 to recover an isometry of $\bar{S}$.